MULTISPECTRUM ANALYSIS OF THE ν_{1} BAND OF HCN: INTENSITIES, BROADENING AND SHIFT COEFFICIENTS

V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; \underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; S. W. SHARPE and R. L. SAMS, Mail Stop K8-88, Battelle Blvd., Pacific Northwest National Laboratory, P. O. Box 999, Richland, WA 99352.

Infrared spectra of HCN in the 3200−3400 cm^{−1} region have been recorded at 0.008 and 0.005 cm^{−1} resolution using two different Fourier transform spectrometers (FTS), the McMath-Pierce FTS at the National Solar Observatory on Kitt Peak and the Bruker IFS 120 HR FTS at the Pacific Northwest National Laboratory. Spectra were recorded with pure HCN at room temperature as well as mixtures of HCN in N_{2} or HCN in air at temperatures from +26^{∘}C to −60^{∘}C. Using our multispectrum technique,^{a} we fit up to 27 spectra simultaneously to determine positions, absolute intensities, self-broadening and self-shift coefficients at room temperature for lines belonging to the H^{12}C^{14}N ν_{1} band centered near 3311 cm^{−1}. In addition, we were able to determine intensities, self-broadening and self-shift coefficients for some lines of the ν_{1} + ν_{2}^{1} − ν_{2}^{1} hot band and several line intensities for H^{13}C^{14}N and H^{12}C^{15}N. The measured line intensities in the ν_{1} band of the main isotope were further analyzed to derive the vibrational band intensity and Herman-Wallis coefficients.^{b} Parameters for broadening and shifts by N_{2}, along with their temperature-dependences, were determined from simultaneous fits to 27 spectra, and we fit 22 spectra simultaneously to obtain the air-broadening and shift parameters. We were able to determine these parameters for both broadening gases for transitions of the ν_{1} band of H^{12}C^{14}N with assignments between P(29) and R(28).^{c} Differences between our results, previous measurements, and the parameters in the current HITRAN database^{d} will be quantified and discussed.

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

bV. Malathy Devi et al., JQSRT, in press (2003).

cC. P. Rinsland et al., JQSRT, in press (2003).

dL. S. Rothman et al., JQSRT, in press (2003).